Line AB↔ passes through the points A(3,−1) and B(1,7).
Which lines are parallel to AB↔?


y−1=2(x+2)
y=−4x+8
y−6=−4(x−3)
y+1=−4(x+1)
y=−4x
y=2x−3
y=4x+7
y=4x

Answer by Boreal(15235)

The slope is (7-(-1)/(1-3)=8/(-2)=-4
The second one is parallel, the third one, the fourth one, and the fifth one. All of them are in the form y=-4x+ some number.